OpenStudy (anonymous):

a box has a volume of 48 cubic feet. if the base of the box is 6 feet by 4 feet what is the height

OpenStudy (anonymous):

volume=area of base * height can you continue?

OpenStudy (anonymous):

(confused)

OpenStudy (anonymous):

yu just should put the values in equation. 48=(area of base)*height area of base=6*4=24 48=24*height does it make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so you find the answer?

OpenStudy (anonymous):

um not really

OpenStudy (anonymous):

48=24*height so height=48/24=2 did you get it?

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

welcome :)